# Federated Variational Inference for Bayesian Mixture Models

### v1

@@ I.5 Number of Batches
$N$ Model ARI Clusters Time (s) Rel Irrel
20000 full 0.865 [0.862, 0.868] 15 [15, 15] 782 [687, 835] 80 [80, 80] 18 [17, 18]
par 0.865 [0.862, 0.868] 15 [15, 15] 207 [187, 228] 80 [80, 80] 17 [16, 17]
FedMerDel 0.850 [0.844, 0.857] 15 [15, 15] 223 [212, 242] 80 [80, 80] 20 [20, 20]
50000 full 0.885 [0.876, 0.885] 10 [10, 10] 1607 [1493, 1767] 75 [75, 75] 23 [22, 24]
par 0.885 [0.876, 0.886] 10 [10, 10] 462 [440, 527] 75 [75, 75] 23 [22, 24]
FedMerDel 0.881 [0.873, 0.883] 10 [10, 10] 630 [563, 672] 75 [75, 75] 25 [25, 25]
J.3 Number of Batches
Detailed results from this simulation study can be found in Table 14 and Figure S5. A Kruskal-Wallis test testing for any significant difference between the ARI for different numbers of batches for $N=100,000$ gives a p-value of 0.125. For $N=200,000$, the p-value is 0.0938. One way to potentially improve inference for the number of clusters for higher values of $K_{\mbox{init}}$ would be to look at higher frequencies of merge/delete moves within each MerDel run.
Table 14: ‘Number of Batches’ simulation results. We report the median and lower/upper quantiles across all 5 ‘shuffles’ for each of the 20 synthetic datasets. Results use the random search for the global merge; there were only marginal differences for the greedy search.
$N$ Batches ARI Clusters
100000 2 0.956 [0.950, 0.960] 10 [10, 10]
100000 4 0.955 [0.951, 0.960] 10 [10, 10]
100000 10 0.955 [0.951, 0.959] 10 [10, 11]
200000 4 0.954 [0.946, 0.959] 10 [10, 10]
200000 8 0.954 [0.946, 0.959] 10 [10, 10]
200000 20 0.953 [0.944, 0.958] 10 [10, 11]
Figure S5: Boxplot showing the distribution of ARI scores across different numbers of batches in the ‘Number of Batches’ simulation.
J.4 Comparing Merge Criteria
In this simulation, we compared methods of selecting candidate clusters for merging as described in Section E.1.1. We compared a correlation measure (‘corr’) and two divergence-based measures (Kullback-Leibler divergence, ‘KL’, and the Bhattacharyya distance, ‘bha’) to completely random selection of candidate clusters (‘rand’).
We ran two simulation studies: one with $N=2000$, $K_{\mbox{true}}=10$, $K_{\mbox{init}}=20$ and $P=50$ (A in Figure S6), and one with $N=2000$, $K_{\mbox{true}}=20$, $K_{\mbox{init}}=40$ and $P=100$ (B in Figure S6). In both cases, we simulated 30 independently generated binary datasets. For each dataset and each model, we ran the model 10 times with a different initialisation each time (giving 40 initialisations per dataset, and 10 per model). For each initialisation, we tested laps = {1, 5}.
Figure S6 illustrates that generally, there was not a significant difference between the choice of methods to select candidate clusters for merging in terms of time and accuracy, although as expected, random choice did tend to take slightly longer - especially for laps = 5 and in the dataset with more clusters. In plot B in Figure S6(b), the median of the ARIs achieved by ‘rand’ is slightly lower than other methods with both laps = 1 and 5. We saw that the proportion of accepted merges was significantly lower than all other methods with random choice in most cases, and that ‘corr’ saw slightly lower merge rates than divergence-based measures. However, the computational complexity of calculating divergences is higher, leading to similar total times for the overall algorithm in Figure S6(a).
@@
Figure S6: Boxplots comparing the distributions of ARI scores, (the logarithm of) total wall-clock time until convergence, and ‘merge rates’. All initialisations for each model, simulated dataset and ‘laps’ value are included, where the boxplots illustrate the median, quantiles and range across all clustering solutions.
J.5 Categorical Data Simulations
J.5.1 Frequency of Merge/Delete Moves
@@
Details of the scenarios for this simulation study with categorical data (analogous to Section I.3) are given in a table below. Both simulations were run without variable selection. $K_{init}$ refers to the initialised value of $K$ in the model. In both scenarios, we generated 20 independent datasets and compared 10 different initialisations for each dataset.
@@
Table 15: Table giving parameters for data generation for the frequency of merge/delete moves with categorical data.
ID $N$ $P$ $K_{init}$ $K_{true}$ $N$ per Categories
Cluster
cat.1 2000 100 20 8 50-800 4
cat.2 4000 100 25 10 200-800 3
Tables comparing wall-clock time, final log-ELBO and ARI of the resulting clustering structures in Simulations cat.1, and cat.2 are shown below (Tables 16, 17).
It is notable that the model with no merge/delete moves (laps = 10000) performed significantly worse in terms of ARI in these simulations, especially when we increased the number of categories in each covariate (ie. in Simulation cat.1 with 4 categories). We also saw in both simulations that the fully variational model never ended with fewer clusters than the initialisation, which could suggest that the lower ARI could be associated with observations remaining in smaller clusters which are subsets of the original clusters. ARI scores only take into account pairwise clustering similarity; one cluster being split into multiple subclusters in another clustering result will result in the two clustering structures sharing a relatively low ARI, due to observations being in different clusters, even though it is a subcluster of the original cluster.
A brief inspection of the resulting clustering structures in Simulation cat.1 (for example) provided further evidence to this effect, where many of the 20 clusters contained more than just a few ‘left-out’ observations. In Simulation cat.1, the very uneven true cluster sizes could have also contributed to this effect. This further supports the usefulness of merge/delete moves in MerDel when applied to categorical data; these moves clearly help to mitigate poor optima and find a number of clusters closer to the truth.
Figure S9 illustrates the change in ELBO versus time for one simulated dataset for both categorical simulations. It is clear in these graphs that MerDel is able to converge at a ‘better’ local optima of the optimisation surface (ie. one with a higher ELBO) compared to the usual variational algorithm.
Figure S7: Scatter plot comparing ARIs achieved by each model across all datasets and initialisations in Simulations cat.1 and cat.2 (labelled 1, 2 respectively). Each point represents one ARI from one MerDel run.
Figure S8: Plot comparing final number of clusters in clustering models in simulations cat.1 and cat.2.
@@
Table 16: Table comparing the mean wall-clock time, mean final log-ELBO, mean ARI and mean number of clusters of the final labelling structure across all runs in Simulation cat.1 with confidence intervals for time and ARI (+- 1.96 x standard deviation).
Laps Time (s) log-ELBO ARI Clusters
0 14.4 [14.2, 14.5] -238524 0.961 [0.951, 0.971] 7.72
1 27.6 [25.6, 29.5] -238378 0.994 [0.993, 0.995] 7.48
2 24.5 [23.1, 25.8] -238356 0.995 [0.994, 0.996] 7.56
5 29.2 [28.5, 29.8] -238352 0.995 [0.994, 0.996] 7.71
10 38.7 [37.8, 39.5] -238513 0.986 [0.984, 0.988] 8.37
10000 79.2 [74.6, 83.8] -241501 0.831 [0.827, 0.835] 20
Table 17: Table comparing the mean wall-clock time, mean final log-ELBO, mean ARI and mean number of clusters of the final labelling structure across all runs in Simulation cat.2 with confidence intervals for time and ARI (+- 1.96 x standard deviation).
Laps Time (s) log-ELBO ARI Clusters
0 34.8 [34.1, 35.4] -360586 0.991 [0.988, 0.995] 10.1
1 47.0 [46.0, 48.1] -360544 0.999 [0.999, 1.00] 10.0
2 49.8 [48.8, 50.7] -360544 0.999 [0.999, 1.00] 10.0
5 66.1 [65.1, 67.0] -360568 0.998 [0.998, 0.999] 10.2
10 91.1 [89.5, 92.8] -360795 0.991 [0.990, 0.992] 12.1
10000 172 [162, 181] -312628 0.939 [0.937, 0.941] 25
(a) Simulation cat.1
(b) Simulation cat.2
Figure S9: Graphs showing the ELBO vs wall-clock time for the algorithm for two simulated datasets across all 20 initialisations for each of the different ”laps” in Simulations cat.1 and cat.2. The mean end time across all initialisations is shown as well as an approximate 95 confidence interval.
J.5.2 Global Merge Simulations
For this simulation with categorical data, we employed the ‘random search’ for the global merge and used the KL divergence to assess similarities between clusters in both the local merge move and the global merge (recall in the binary simulation, we used correlation). The three scenarios had simulated data with $N=$ 20,000, 50,000, 100,000, with $K=$ 20, 10 and 12 true clusters split into 4, 5 and 5 equally sized batches respectively. We set $P=100$ covariates and 3 categories in each covariate in all cases. We generated 5 simulated datasets and had 10 different ‘shuffles’ of the data in each scenario, as well as running MerDel and parallelised MerDel on the full data 10 times, as with the binary simulations.
Results showed extremely good performance with almost perfect clustering structures whether or not we employed a global merge or not in all cases. There was a clear speed up for the global merge model and for parallelised MerDel, and as we saw with binary data, as the data size increased, we saw increased gains from using FedMerDel as opposed to parallelising MerDel stepwise. Note the higher merge times and total times with $N=20,000$ were likely a result of the increased number of true clusters and initialised clusters in each MerDel run (30). In the later two simulations, we initialised with 20 clusters.
Table 18: ‘Global Merge Simulations’ with categorical data results. We report the median and lower/upper quantiles across all 10 ‘shuffles’/initialisations and all 5 synthetic datasets.
@@
$N$ Model ARI Clusters Time (s) Global Merge Time (s)
20000 full 0.999 [0.999, 0.999] 20 [20, 20] 387 [373, 418]
par 0.999 [0.999, 0.999] 20 [20, 21] 146 [128, 157]
FedMerDel 0.998 [0.99, 0.998] 20 [20, 20] 124 [118, 131] 3.91 [3.82, 4.14]
50000 full 0.999 [0.999, 1.00] 10 [10, 10] 374 [373, 375]
par 0.999 [0.999, 1.00] 10 [10, 10] 127 [125, 128]
FedMerDel 0.999 [0.999, 1.00] 10 [10, 10] 83.5 [80.7, 84.1] 1.29 [1.28, 1.34]
100000 full 1.00 [0.999, 1.00] 12 [12, 12] 946 [857, 1050]
par 1.00 [0.999, 1.00] 12 [12, 12] 293 [276, 339]
FedMerDel 0.999 [0.990, 1.00] 12 [12, 13] 243 [219, 280] 2.57 [2.49, 2.64]
Appendix K MNIST - Additional Results
K.1 Supplementary Figures
(a) Original
(b) Dimension-reduced
(c) Binarised
@@
Figure S10: Example of the first 10 MNIST test-set digits visualised in their original image format (28x28), the same digits with the dimensionality reduction applied (16x16), and the same dimension reduced digits converted to a binary format.
(a) 4
(b) 7
(c) 9
@@
Figure S11: 10 randomly selected digits of 4, 7 and 9 from the MNIST dataset, showing similarities in their shapes.
@@
Figure S12: A heatmap showing the correspondence between clusters and true numbers in the clustering model with the best ELBO for the MNIST data. A darker cell colour in entry (i, j) indicates a higher percentage of samples from number i are in the given cluster j. The exact number corresponds to the percentage of cluster j made up of the given number.
(a) Sorted by true digits
(b) Sorted by cluster
Figure S13: Heatmaps showing the first 2000 digits in our pre-processed MNIST dataset, first sorted by true digit and then sorted by cluster for the clustering model with the best ELBO. Differences between and within digits can be seen.
(a) Columns in default order
(b) Columns shuffled
Figure S14: Heatmaps showing only 0’s and 1’s in the first 10000 digits in our pre-processed MNIST dataset, sorted by cluster within both digits. Shuffling the columns allows the differences between separate clusters to be clearer.
K.2 Supplementary Results - Classification Rates
@@
  • FedMerDel_rand: Our novel federated learning model, FedMerDel, analysed in the main paper. This included 6 equally sized batches of data with 20 maximum clusters in each batch. A random search was used for the global merge. All models took between 77-126 mins.
  • FedMerDel_greedy: FedMerDel, but with a greedy search used for the global merge. The same batches were used as with FedMerDel_random above. All models took between 77-127 minutes.
  • FedMerDel_30r: FedMerDel with 6 equally sized batches of data, but now with 30 maximum clusters in each batch. These models had slightly better classification rates, but came at a high cost of computational time with 60 extra clusters being processed, and reduced scope for interpretation of clusters. Resulting models had between 52-70 clusters. More clusters than 30 per batch could be used if an even finer and robust clustering structure is desired, but this was not the aim of our analysis, where we focused on finding a broad clustering structure with a large dataset to validate the scalability of our method. This model used a random search.
@@
  • kmeans_xx_yy: In this model, in order to implement scalable clustering methods for continuous data to compare to, we projected the data to a lower dimension with Multiple Correspondence Analysis (MCA), a counterpart to principal component analysis for categorical data. We then implement the k-means algorithm on this lower-dimesional data. xx represents the number of MCA dimensions kept in the model, which is 5 (explaining 20% of variance), 10 (explaining 30% of variance) or 17 (explaining 40% of variance). yy represents the number for $k$, the number of clusters, where we test 20 and 35. In all cases, we use 20 maximum iterations for k-means, and 20 different initialisations, and use the clustering with the best total within-cluster sum of squares. We saw that this method is quick (taking no longer than a minute) and generally performs well especially as we increase the number of dimensions and clusters, although does not separate numbers such as 2, 3 and 6 as well as our models.
  • EM_yy: We applied an EM algorithm for frequentist model estimation to fit a finite mixture model to the MNIST binary data via the R package flexmix (Leisch, 2004) and used the BIC (Bayesian Information Criterion) for model selection across a given number of initialisations. yy represents the number of clusters in the finite mixture model, and we test 20 (10.2 mins) and 35 (55.4 mins). We use 10 different initialisations, and the model returned is the maximum likelihood solution.
  • MCA_EM_yy: We also applied an EM algorithm to fit a Gaussian mixture model to MCA-transformed MNIST data. For this, we used the R package mclust which also uses the BIC to select the optimal model between 14 models with different shapes, volumes, and orientations of covariances. yy represents the number of clusters in the finite mixture model, and we test 20 (4.45 mins) and 35 (9.77 mins).
Generally, FedMerDel performed similarly well to comparitor methods, where 0’s, 1’s, 2’s and 6’s saw higher classification rates with methods such as EM.
Table 19: Table comparing classification rates (%) for each digit across different clustering models. The model analysed further in the paper and supplement was the 4th FedMerDel_rand run as the model with the highest ELBO of all the FedMerDel_rand and FedMerDel_greedy runs. Classification rate is defined as the percentage of digits classified into a cluster primarily consisting of that digit.
@@
Figure S15: 3 randomly selected observations and 2 generated digits (from the variational distribution) from each cluster in the clustering model with the best ELBO for the MNIST data.
Appendix L EHR Data - Additional Information
@@
EHR data were obtained from The Health Improvement Network database. The start date for our study was 01 January 2010 and end date was 01 January 2017. The dataset comprised primary care health data for 289,821 individuals over 80 years old who had at least two of the long term health conditions considered in our analysis. The full list of health conditions, and their coding, is provided in Table 20. 64.8% of individuals contributing to the analysis dataset were female, 35.2% were male.
L.1 Greedy Search Global Merge
@@
In Figure 3 of the main manuscript, we illustrated the global clusters identified using the “random search” approach to perform the global merge. In Figure S16 below, we provide the corresponding results obtained using the “greedy search” approach.
@@
Figure S16: Results of global EHR clustering using the greedy search approach to perform the global merge. Columns are health conditions, and rows correspond to clusters. Row labels (shown on the right) indicate the number of individuals in each cluster. Summary names for each of the clusters are provided on the left of the plot. To improve visualisation, only the most prevalent health conditions are shown.

### v2

@@ H.2.4 Global Merge with Variable Selection
H.3 Number of Batches
Results from this simulation study are visualised in Figure 9. A Kruskal-Wallis test testing for any significant difference between the ARI for different numbers of batches for $N=100,000$ gives a p-value of 0.125. For $N=200,000$, the p-value is 0.0938. One way to potentially improve inference for the number of clusters for higher values of $K_{\mbox{init}}$ would be to look at higher frequencies of merge/delete moves.
Figure 9: Boxplot showing the distribution of ARIs across different numbers of batches in the ‘Number of Batches’ simulation.
H.4 Comparing Merge Criteria
In this simulation, we compared methods for selecting candidate clusters for merging (Section C.1.1): correlation (corr’), Kullback-Leibler divergence (KL’), Bhattacharyya distance (bha’), and random selection (rand’).
We ran two studies with $N=2000$ and different complexity: (A) $K_{\mbox{true}}=10$, $K_{\mbox{init}}=20$, $P=50$; (B) $K_{\mbox{true}}=20$, $K_{\mbox{init}}=40$, $P=100$ (Figure 10). In both cases, we simulated 30 independently generated binary datasets. For each dataset, we tested 4 models with 10 initializations each and laps = 1, 5.
Figure 10 shows no significant differences between methods in time and accuracy, though random selection took slightly longer (especially with laps = 5 and more clusters) and achieved slightly lower median ARIs in study B. Random selection had significantly lower merge acceptance rates, while correlation had lower rates than divergence measures. However, higher computational complexity of divergences resulted in similar total algorithm times.
@@ H.4 Comparing Merge Criteria
Figure 10: Boxplots comparing the distributions of ARI scores, (the logarithm of) total wall-clock time until convergence, and ‘merge rates’. All initialisations for each model, simulated dataset, and ‘laps’ value are included, where the boxplots illustrate the median, quantiles, and range across all clustering solutions.
H.5 Categorical Data Simulations
H.5.1 Frequency of Merge/Delete Moves
@@
Details of the scenarios for this simulation study with categorical data (analogous to Section G.3) are given in a table below. Both simulations were run without variable selection. In both scenarios, we generated 20 independent datasets and compared 10 different initialisations. Table LABEL:catsim_combined shows results; the model with no merge/delete moves (laps = 10000) performed markedly worse in ARI.
[…]
simFedMerDeltablecat ID $\mathbf{N}$ $\mathbf{P}$ $\mathbf{K_{init}}$ $\mathbf{K_{true}}$ $\mathbf{N}$ per Categories Cluster cat.1 2000 100 20 8 50-800 4 cat.2 4000 100 25 10 200-800 3
@@
Table 11: Parameters for data generation for frequency of merge/delete moves with categorical data.
[…]
catsim_combined Simulation Laps Time (s) log-ELBO ARI Clusters cat.1 0 14.4 [14.2, 14.5] -238524 0.961 [0.951, 0.971] 7.72 1 27.6 [25.6, 29.5] -238378 0.994 [0.993, 0.995] 7.48 2 24.5 [23.1, 25.8] -238356 0.995 [0.994, 0.996] 7.56 5 29.2 [28.5, 29.8] -238352 0.995 [0.994, 0.996] 7.71 10 38.7 [37.8, 39.5] -238513 0.986 [0.984, 0.988] 8.37 10000 79.2 [74.6, 83.8] -241501 0.831 [0.827, 0.835] 20 cat.2 0 34.8 [34.1, 35.4] -360586 0.991 [0.988, 0.995] 10.1 1 47.0 [46.0, 48.1] -360544 0.999 [0.999, 1.00] 10.0 2 49.8 [48.8, 50.7] -360544 0.999 [0.999, 1.00] 10.0 5 66.1 [65.1, 67.0] -360568 0.998 [0.998, 0.999] 10.2 10 91.1 [89.5, 92.8] -360795 0.991 [0.990, 0.992] 12.1 10000 172 [162, 181] -312628 0.939 [0.937, 0.941] 25
@@
Table 12: Comparisons of mean wall-clock time, mean final log-ELBO, mean ARI and mean number of clusters of final labelling structure across all runs in Simulations cat.1, cat.2 with confidence intervals for time and ARI (+- 1.96 x standard deviation).
H.5.2 Global Merge Simulations
We employed random search for global merging using KL divergence to assess cluster similarities. Three scenarios used simulated data with $N=$ 20,000, 50,000, 100,000 observations, $K=$ 20, 10, 12 true clusters, and 4, 5, 5 equally-sized batches respectively. All scenarios had $P=100$ covariates with 3 categories each. With 5 datasets and 10 runs per model, results showed near-perfect clustering. FedMerDel and parallelised MerDel improved speed, with FedMerDel showing increasing efficiency gains as $N$ increased.
[…]
@@ H.5.2 Global Merge Simulations
globalmergeresultscat $\mathbf{N}$ Model ARI Clusters Time (s) Global Merge Time (s) 20000 full 0.999 [0.999, 0.999] 20 [20, 20] 387 [373, 418] par 0.999 [0.999, 0.999] 20 [20, 21] 146 [128, 157] FedMerDel 0.998 [0.99, 0.998] 20 [20, 20] 124 [118, 131] 3.91 [3.82, 4.14] 50000 full 0.999 [0.999, 1.00] 10 [10, 10] 374 [373, 375] par 0.999 [0.999, 1.00] 10 [10, 10] 127 [125, 128] FedMerDel 0.999 [0.999, 1.00] 10 [10, 10] 83.5 [80.7, 84.1] 1.29 [1.28, 1.34] 100000 full 1.00 [0.999, 1.00] 12 [12, 12] 946 [857, 1050] par 1.00 [0.999, 1.00] 12 [12, 12] 293 [276, 339] FedMerDel 0.999 [0.990, 1.00] 12 [12, 13] 243 [219, 280] 2.57 [2.49, 2.64]
Table 13: ‘Global Merge Simulations’ with categorical data results. We report the median and lower/upper quantiles across all 10 ‘shuffles’/initialisations and all 5 synthetic datasets. Note the higher merge times and total times with $N=20,000$ is likely due to more initialised clusters (30 vs. 20 in the larger NN settings).
Appendix I MNIST Study
I.1 Simulation Set-Up
We applied FedMerDel to the MNIST collection of $N=60000$ images of handwritten digits 0-9 (MNIST). As in JMLR:v17:11-392, we downsampled each image to $16\times 16$ pixels using bilinear interpolation via TensorFlow, and represented this as a 256-dimensional binary vector. Furthermore, to reduce computation on irrelevant pixels (e.g. those in corners of the image) we removed all pixels with fewer than 100 non-zero values across the whole dataset, leaving 176 covariates . We split this into 6 batches of $N=10000$ and set $K=20$ maximum clusters per batch. True labels for the MNIST digits were discarded.
I.2 Results
We ran FedMerDel on 5 different ‘shuffles’ of the data into batches with $K=20$ maximum clusters in each batch, and our resulting models after a global merge consisted of 27-33 clusters. The observed variance in clustering outcomes reflects the genuine difficulty of this unsupervised task - handwritten digits exhibit substantial within-class variation when represented as binary vectors, and this variance is comparable to other unsupervised clustering methods on MNIST. It is unrealistic to expect to find 10 clusters exactly due to the variation of handwritten digits within numbers. The increased number of clusters was not an artefact of the global merging; a run of MerDel on the MNIST test set of $N=10,000$ digits with 35 maximum clusters found between 33-35 clusters where most merges/deletes were rejected. Other unsupervised clustering methods also found a high number of clusters for MNIST (hughes_memoized_2013; Ni2020).
Figure 13 depicts the distribution of true number labels within each cluster, and demonstrates subclustering within digits in the ‘best’ model out of our 5 clustering structures. This is chosen as the model with the highest ELBO, an approach used in other variational models (Ueda2002; hughes_memoized_2013). FedMerDel achieved more consistent performance on certain well-separated digits: for example, 95.95% of 1’s and 92.46% of 6’s were classified into clusters predominantly made up of that digit, while digits with similar shapes (4, 7, 9) were more challenging to distinguish. These numbers have quite similar shapes and have a variety of shapes (Figure 12). We expected imperfect performance due to the reduction of the image to a 1D vector; every pixel is treated as independent of one another, and we lose all information from the 2D images pertaining to correlation between nearby pixels. Further classification rates can be found in Appendix I.3, as well as a comparison to other scalable clustering methods.
This analysis serves as a supplementary benchmark on real-world data with ground truth labels, while our primary application domain remains EHR data where our simulated experiments demonstrate much lower sensitivity to batch partitioning.
We can use the posterior distribution of our model to probabilistically produce new digits from each cluster, showing the generative potential in variational Bayes; see Figure 14.
@@ I.2 Results
Figure 11: Example of the first 10 MNIST test-set digits visualised in their original image format (28x28), the same digits with the dimensionality reduction applied (16x16), and the same dimension reduced digits converted to a binary format.
@@
Figure 12: 10 randomly selected digits of 4, 7 and 9 from the MNIST dataset, showing similarities in their shapes.
@@
Figure 13: A heatmap showing the correspondence between clusters and true numbers in the clustering model with the best ELBO for the MNIST data. A darker cell colour in entry (i, j) indicates a higher percentage of samples from number i are in the given cluster j. The exact number corresponds to the percentage of cluster j made up of the given number.
I.3 Supplementary Results - Classification Rates
@@
  • FedMerDel_rand: FedMerDel, with 6 equally sized batches of data with 20 maximum clusters in each batch. A random search was used for the global merge. All models took between 77-126 mins.
  • FedMerDel_greedy: FedMerDel, but with a greedy search used for the global merge. The same batches were used as with FedMerDel_rand above. All models took between 77-127 minutes.
  • FedMerDel_30r: FedMerDel with 6 equally sized batches of data, but now with 30 maximum clusters in each batch. These models came with high computational time with 60 extra clusters being processed, and reduced scope for interpretation of clusters. Resulting models had between 52-70 clusters. This model used a random search.
@@
  • kmeans_xx_yy: Data was projected to a lower dimension with Multiple Correspondence Analysis (MCA). We then implemented the k-means algorithm on this lower-dimensional data. xx represents the number of MCA dimensions kept in the model: 10 (explaining 30% of variance) or 17 (explaining 40% of variance). yy represents the number for $k$, the number of clusters, where we test 20 and 35. All cases use 20 maximum iterations for k-means, and 20 different initialisations, and present the clustering with the best total within-cluster sum of squares. This method was quick (taking no longer than a minute) and generally performed well, but did not separate numbers such as 2, 3 and 6 as well as FedMerDel.
  • EM_yy: We applied an EM algorithm for frequentist model estimation to fit a finite mixture model to the MNIST binary data via the R package flexmix (Leisch2004) and used the BIC (Bayesian Information Criterion) for model selection across 10 initialisations. yy represents the number of clusters in the finite mixture model, and we test 20 (10.2 mins) and 35 (55.4 mins).
Generally, FedMerDel performed similarly well to comparator methods.
Table 14: Table comparing classification rates (%) for each digit across different clustering models. The model analysed further was the 4th FedMerDel_rand run as the model with the highest ELBO of all the FedMerDel_rand and FedMerDel_greedy runs. Classification rate is defined as the % of digits classified into a cluster primarily consisting of that digit.
@@
Figure 14: 3 randomly selected observations and 2 generated digits (from the variational distribution) from each cluster in the clustering model with the best ELBO for the MNIST data.
Appendix J EHR Data - Additional Information
@@
EHR data were obtained from The Health Improvement Network database. The start date for our study was 01 January 2010 and end date was 01 January 2017. The dataset comprised primary care health data for 289,821 individuals over 80 years old who had at least two of the long term health conditions considered in our analysis. The full list of health conditions, and their coding, is provided in Table 17. Prevalence of the conditions in Figure LABEL:THINsummary is in Table 16. 64.8% of individuals contributing to the analysis dataset were female, 35.2% were male.
J.1 Characterisation of Random Search Clusters
Note the mean mortality rate across the full dataset was 0.48, and the mean number of conditions per person was 6.19. Mortality rate is defined as the number of people with a registered death date on or before 31/12/2020.
Table 15: Characterising the clusters seen in Figure LABEL:THINsummary by the mean number of conditions and mortality rate.
Cluster Mean No. of Conditions Mortality
(i) cancer 6.90 0.49
(ii) af & arrhythmia 8.48 0.57
(iii) pvd & aortic aneurysm 8.77 0.61
(iv) stroke 8.46 0.54
(v) af & arrhythmia & stroke 11.3 0.61
(vi) stroke & blindness 12.6 0.67
(vii) blindness 8.69 0.53
(viii) dementia 6.14 0.61
(ix) asthma & copd 6.30 0.54
(x) ‘generally multimorbid’ 4.57 0.39
(xi) ‘generally multimorbid’ 9.40 0.41
(xii) ‘generally multimorbid’ 7.85 0.55
Table 16: Table providing prevalence as a percentage for the conditions in Figure LABEL:THINsummary. The most prevalent condition not included in this table is gout, at a prevalence of 7.63%, which is a condition often found as a comorbidity, but it is not expected that this should form a separate cluster. Other conditions are less common.
Coding Prevalence Coding Prevalence Coding Prevalence
cancerall 13.8 tia_stroke 18.5 anxiety 9.38
ca_skin 12.6 eczema 16.0 dementia 13.1
arrythmia 18.3 deaf 27.0 oa 39.3
af 16.3 blind 5.94 osteoporosis 14.4
hypertension 63.2 cataract 34.9 ckd 33.2
hf 10.1 glaucoma 9.77 asthma 12.2
ihd 26.8 amd 8.47 copd 8.20
pvd 7.64 retinopathy 8.08 hypothyroid 13.1
aortic_aneurysm 1.84 diverticuli 14.6 t2dm 14.7
tia 9.55 depression 13.6 bph 9.43
nos_stroke 10.4
J.2 Heterogeneous EHR Simulations
Figure 15: Results of global EHR clustering where individuals with and without dementia were separated into different clusters. 15 total similarly sized batches were randomly generated after the data was divided. Random merge was used. Similar clusters are seen, but split by those with and without dementia as expected. Notably, an asthma/COPD cluster emerges in patients without dementia, but this is not seen for patients with dementia. Similar mortality/mean number of conditions as in Appendix J.1 could be found to infer clinical risk based on having dementia as an additional comorbidity or not. To improve visualisation, only the most prevalent health conditions are shown.
Figure 16: Results of global EHR clustering where a symmetric 15-dimensional Dirichlet(1) distribution was used to simulate heterogeneity. This was done by taking the 12 clusters from the analysis in the main manuscript, and assigning observations from each cluster into 15 new batches according to the Dirichlet distribution. Random merge was used. Core clusters are still clearly retained (eg. cancer, respiratory disorders, stroke only, mixed stroke + AF/arrhythmia) while smaller clusters emerge, likely driven by shard specific overrepresentation. For example, a small multimorbid cluster emerges with no osteoarthritis. To improve visualisation, only the most prevalent health conditions are shown.
J.3 Greedy Search Global Merge
@@
In Figure 3 of the main manuscript, we illustrated the global clusters identified using the “random search” approach to perform the global merge. In Figure 17 below, we provide the corresponding results obtained using the “greedy search” approach.
@@
Figure 17: Results of global EHR clustering using the greedy search approach to perform the global merge. Columns are health conditions, and rows correspond to clusters. Row labels (shown on the right) indicate the number of individuals in each cluster. Summary names for each of the clusters are provided on the left of the plot. To improve visualisation, only the most prevalent health conditions are shown.
J.4 Full List of Health Conditions